OpenStudy (anonymous):

I don't know how to do this equation.. (5x^2 + x – 4)(x + 2)

OpenStudy (anonymous):

Make the factors of first parenthesis.. See tell me the two numbers with sum = 1 and product = -20??

OpenStudy (anonymous):

I don't know how to do that.

OpenStudy (anonymous):

See, try with 4 and 5.. we need -20.. in what way I multiply 4 and 5 to get -20?

OpenStudy (anonymous):

Negative + positive?

OpenStudy (anonymous):

Which one should be negative and which one should be positive?

OpenStudy (anonymous):

It doesn't matter. Either one works.. (-4)(5)=-20 and (-5)(4)=-20

OpenStudy (anonymous):

But if I say that their sum should be 1 then which sign you will apply to which?

OpenStudy (anonymous):

I don't know.

OpenStudy (anonymous):

See, you told we can take -4 and 5 or -5 and 4 just add them to get 1. which will it adds up to 1? -5 + 4 Or, -4 + 5 ??

OpenStudy (anonymous):

-4+5

OpenStudy (anonymous):

Yes now you got it.. or if I multiply x to both -4 and 5 what you get?? x(-4 + 5) distribute x inside..

OpenStudy (anonymous):

-4x+5x

OpenStudy (anonymous):

Yes so you agree? 1x = (-4 + 5)x ??

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

Now tell me the expression in first parenthesis?

OpenStudy (anonymous):

Um.. what expression?

OpenStudy (anonymous):

See in the question what is the expression in first bracket? it is quadratic for which we are solving for..

OpenStudy (anonymous):

5x^2 ?

OpenStudy (anonymous):

tell me the whole..

OpenStudy (anonymous):

5x^2 + x - 4

OpenStudy (anonymous):

Yes.. now I make you agree that 1x = (5 - 4)x Yes or not?

OpenStudy (anonymous):

Is x = (5-4)x right??

OpenStudy (anonymous):

Yes?

OpenStudy (anonymous):

Then replace the middle term of the expression: 5x^2 + x - 20 Here you can replace middle term x with (5-4)x can you do it?

OpenStudy (anonymous):

5x^2 + (5-4) - 20 But how did you replace the - 4 with the 20?

OpenStudy (anonymous):

no no.. See we are to replace middle term only using the two numbers we got.. See: \[5x^2 + x -20 = 5x^2 + 5x - 4x -20\] This step are you getting?? I just repace x by (5-4)x and then distribute x to get 5x - 4x Getting or not?

OpenStudy (anonymous):

Yes, I understand.

OpenStudy (anonymous):

Now take common that can you take out..

OpenStudy (anonymous):

See: from: \[5x^2 + 5x\] What can be taken out of it as common?

OpenStudy (anonymous):

The 5x?

OpenStudy (anonymous):

Yes.. so you have: \[5x(x + 1) -4x - 20\] Now from the next part what can be taken out?

OpenStudy (anonymous):

How did you get 5x(x+1)-4-20 ?

OpenStudy (anonymous):

After taking 5x as common tell me what you are left with?

OpenStudy (anonymous):